OpenStudy (loser66):

Find z such that tanz = i/2 Please, help

OpenStudy (loser66):

My attempt: \(tan z = \dfrac{sin z}{cos z}=\dfrac{e^{iz}-e^{-iz}}{2i} *\dfrac{2}{e^{iz}+e^{-iz}}=\dfrac{i}{2}\)

OpenStudy (loser66):

hey, faking girl, help me out. hehehe

OpenStudy (dan815):